In fourteen years time a mother will be twice as old as her son.Four years ago the sum of their ages was 30 years.Find how old the mother was when the son was born.
Answer by mananth(16946)

In fourteen years time a mother will be twice as old as her son.Four years ago the sum of their ages was 30 years.Find how old the mother was when the son was born.
let mother be x years and son be y years now
After 14 years
(x+14)=2(y+14)
x+14=2y+28
x-2y=14.......................(1)
4 years ago
(x-4) +(y-4) =30
x+Y-8=30
x+y = 38.......................(2)
subtract (2) from (1)
-3y=-24
y=8
Therefore x=30 because x+y=38
Son was born 8 years ago
so mother was 30-8 = 22 years then